Can you please help me solve this equation for the variable specified? thank you
9a-2b=c+4a, for a

Answer by rfer(16322)

9a-2b=c+4a
subtract 4a
5a-2b=c
add 2b
5a=c+2b
divide by 5
a=(c+2b)/5